# Measuring the Spin of the Galactic Center Supermassive Black Hole with Two Pulsars

### v1

@@ Abstract
Measuring the BH spin is crucial for gravity tests. Determining the spin of Sgr A* based on the timing of a single pulsar suffers from parameter degeneracy at the leading order Liu et al. (2012); Zhang and Saha (2017). The measurable secular changes of the pulsar orbit in a short time, caused by the frame-dragging effect Barker and O’Connell (1975), are the linear rate of the periastron precession, $\dot{\omega}$, and the linear change of the projected semimajor axis of the pulsar orbit, $\dot{x}$. However, these two quantities cannot fully determine the BH spin which has three components. To break the degeneracy, one needs at least one accurate measurement of higher-order time derivatives such as $\ddot{\omega}$ or $\ddot{x}$ Liu et al. (2012), which in general requires a much longer observation time span or a pulsar in an orbit with very short orbital period whose relativistic effects are large.
There are still two possible ways to break the leading-order degeneracy and provide a better spin determination without requiring a pulsar in a rare orbital situation. One is to combine the measurement of the proper motion of the pulsar from astrometric observations Zhang and Saha (2017). One can infer the secular change in the longitude of the ascending node, $\dot{\Omega}$—which is not observable in timing observation—from the proper motion of the pulsar and provide an additional constraint on the spin of the central BH. However, an expected astrometric accuracy in the order of $10\,\mu{\rm as}$ Fomalont and Reid (2004) is still not enough and contributes little when combining with the timing data Zhang and Saha (2017). The other possibility is to combine the observation of another pulsar with a different orbital inclination $i$, as the leading-order degeneracy mainly depends on $i$ Zhang and Saha (2017). Considering a potentially large pulsar population in the GC, finding two or more pulsars orbiting around Sgr A* in the future may provide us a more likely way in measuring the spin of Sgr A*.
@@
Combining two pulsars.—When the changes in the Keplerian parameters of a pulsar orbit are slow over the observation time span, the direct way to characterize secular effects is to fit the timing data with timing parameters varying linearly in time, such as $\omega=\omega_{0}+\dot{\omega}\,t$ and $x=x_{0}+\dot{x}\,t$ Damour and Taylor (1992). In a pulsar-SMBH system, the leading-order derivatives will be measured soon after the start of observation and have a rather high measurement precision. Here we use the two combinations introduced by Liu et al. (2012) for the leading-order derivatives and other timing observables, which are $\mathcal{X}\equiv-\dot{x}s_{i}^{2}\big{(}x\hat{\Omega}\big{)}^{-1}$ and $\mathcal{W}\equiv(\dot{\omega}-\dot{\omega}_{M})s_{i}^{2}\hat{\Omega}^{-1}$, where $s_{i}=\sin i$, $\hat{\Omega}=4\pi\beta_{O}^{3}/P_{b}(1-e^{2})^{3/2}$ and $\dot{\omega}_{M}=6\pi\beta_{O}^{2}/P_{b}(1-e^{2})$, with $\beta_{O}=\big{(}2\pi GM/c^{3}P_{b}\big{)}^{1/3}$ the orbital velocity parameter introduced by Damour and Taylor (1992). Note that $s_{i}$ and $M$ in these combinations can be measured precisely from the Shapiro delay Shapiro (1964) that is large enough here even for face-on orbits Liu et al. (2012); Hu et al. (2023).
@@
To combine the timing results of two pulsars and determine the spin of the central BH, one has to translate the measurement constraints of both pulsars in a common parameter space. In this case, the natural choice is the $\chi$-$\chi_{\lambda}$ plane, where $\chi_{\lambda}$ is the projection of the dimensionless spin $\chi$ of the BH in the line of sight direction. For each pulsar, timing observations give one constraint represented by a hyperbola,
$\frac{\mathcal{X}^{2}}{s_{i}^{2}(1-s_{i}^{2})}+\frac{\mathcal{W}^{2}}{s_{i}^{2}(1+3s_{i}^{2})}=\chi^{2}-\frac{1+3s_{i}^{2}}{(1-3s_{i}^{2})^{2}}\left(\chi_{\lambda}+\frac{3\mathcal{W}c_{i}}{1+3s_{i}^{2}}\right)^{2}\,,$ (1)
@@
where $c_{i}=\cos i$. If one observes two or more pulsars, all these curves in the $\chi$-$\chi_{\lambda}$ plane should intersect at one point that gives $\chi$ and $\chi_{\lambda}$.
@@
Figure 1: Illustration of the determination of the spin of Sgr A* with measurements of two pulsars in the $\chi$-$\chi_{\lambda}$ plane. The blue square marks the assumed true value of the BH spin. The solid curves show the constraints from two pulsars, while the dashed lines correspond to the $i\leftrightarrow\pi-i$ ambiguity. Shaded region is excluded by the condition $\left|\chi_{\lambda}\right|\leq\chi$.
@@
Figure 1 shows an illustration of using the timing of two pulsars to determine the spin of Sgr A* in the $\chi$-$\chi_{\lambda}$ plane. The system parameters are randomly chosen to avoid the accidental degeneracy which is discussed later. For each pulsar, the measurements of $({\cal X},{\cal W})$ give a hyperbola shown by the solid curve, and the $i\leftrightarrow\pi-i$ ambiguity gives another hyperbola shown by the dashed line. In general, there can exist at most eight different solutions for $(\chi,\chi_{\lambda})$. A rough measurement of any second-order derivatives, for example, the $\ddot{\omega}$ for one of the pulsar, as shown by the vertical red line, or timing of a third pulsar can reduce this degeneracy into two solutions whose $\chi$ has the same value and $\chi_{\lambda}$ has opposite signs.
@@
The methodology described above can be easily extended for combining more pulsars. For each additional pulsar, it provides an independent test of GR as all these curves should intersect at one point within their measurement uncertainties. On the other hand, if one assumes that GR is correct, timing more pulsars can be a unique probe for us to explore and model the (stellar and gas) mass distribution near the GC Merritt et al. (2010); Liu et al. (2012); Hu et al. (2023). As only the leading-order time derivatives and a rough value of a higher-derivative parameter are used, one is expected to obtain a high-precision measurement after a short time observing two pulsars.
@@
The full parameter set, $\bm{\Theta}=\bm{\Theta}_{\rm BH}\cup\bm{\Theta}_{\rm PSR}$, in the timing model for a single pulsar includes the parameters of the SMBH, $\bm{\Theta}_{\rm BH}=\{M,\chi,q,\lambda,\eta\}$, where $\lambda$ and $\eta$ describe the direction of the spin, and the parameters of the pulsar, $\bm{\Theta}_{\rm PSR}=\{P_{b},e,f,i,\omega,N_{0},\nu,\dot{\nu}\}$, where $f$ is the initial orbital phase and $\{N_{0},\nu,\dot{\nu}\}$ relate the pulsar’s rotation number $N$ and pulsar’s proper time $T$ via $N(T)=N_{0}+\nu T+\dot{\nu}T^{2}/2$. We have set the longitude of the ascending node of the pulsar orbit, $\Omega$, to be zero as it is not an observable for timing observation. When combining the timing of two or more pulsars, one should have one set of $\Theta_{\rm PSR}$ for each pulsar while $\Theta_{\rm BH}$ is common for different pulsars. In addition, the difference between the longitudes of the ascending node, $\Delta\Omega=\Omega_{2}-\Omega_{1}$, should also be included in the timing model Ransom et al. (2014). As we can have an overall rotation of the system around the line of sight direction, we fix $\Omega_{1}=0$ in the following discussions and we denote the pulsar with a smaller orbital period as PSR1 or the inner pulsar.
We use the Fisher matrix method for parameter estimation Damour and Deruelle (1986). The log-likelihood function $\mathcal{L}$ in the covariance matrix $C_{\alpha\beta}=\big{(}\partial^{2}\mathcal{L}/\partial\Theta^{\alpha}\partial\Theta^{\beta}\big{)}^{-1}$ is simply the sum from two pulsars, $\mathcal{L}=\mathcal{L}_{1}+\mathcal{L}_{2}$, and
$\mathcal{L}_{i}=\frac{1}{2\nu_{i}^{2}}\sum_{a=1}^{N^{(i)}_{\rm TOA}}\frac{\left[N^{(i)}_{a}(\bm{\Theta})-N^{(i)}_{a}(\bar{\bm{\Theta}})\right]^{2}}{\left(\sigma^{(i)}_{\rm TOA}\right)^{2}}\,,\ \ i=1,2\,$ (2)
where $N^{(i)}_{a}(\Theta)=N\Big{(}\Theta;t_{a}^{(i)\,{\rm TOA}}\Big{)}$ is the pulsar rotation number calculated by the timing model corresponding to the $a$-th pulse arrived at $t_{a}^{(i)\,{\rm TOA}}$. $\bar{\bm{\Theta}}$ denotes the true system parameters and $\sigma_{\rm TOA}^{(i)}$ is the timing precision for pulsar $i$. As we do not consider the interaction between the two pulsars, the properties of the Fisher matrix method and the timing model allow one to construct the full Fisher matrix $F\equiv C^{-1}$ though the combination of the Fisher matrices $F_{i}$ of each single pulsar. This simplification is important for combining more pulsars.
@@
We assume a timing precision $\sigma_{\rm TOA}=1\,{\rm ms}$ and an observation time span $T_{\rm obs}=5\,{\rm yr}$ for both pulsars, which are realistic for future instruments such as the Square Kilometre Array (SKA) and the next-generation Very Large Array (ngVLA) Weltman et al. (2020); Bower et al. (2018). In the upper panel of Fig. 2 we show the measurement precision of the spin of Sgr A* from combining the timing of two pulsars (solid curves) compared to the results obtained from the inner pulsar only (dashed line). Here we fix the orbital period of the inner pulsar to be $P_{b1}=2.3\,{\rm yr}$ and orbital eccentricity $e_{1}=0.8$ while varying the orbital period $P_{b2}$ and eccentricity $e_{2}$ of the second pulsar. The lower panel of this figure shows the improvement factor $F_{\rm imp}=\delta\chi_{\rm single}/\delta\chi_{\rm combine}$. For this case, the inner pulsar with an orbital period $P_{b1}=2.3\,{\rm yr}$ can only constrain $\delta\chi/\chi$ to $\sim 15\%$. Combining a second pulsar even with a large orbital period can provide an improvement factor of $\sim 10$ because of the breaking of the leading-order degeneracy. The improvement factor from combining a second pulsar will be smaller if the inner pulsar has a tighter orbit for which itself can break the leading order degeneracy. Another case study of $P_{b1}=0.5\,{\rm yr}$ shows an improvement factor of $2\sim 3$ when combined with a second pulsar whose orbital period $P_{b2}\lesssim 3\,{\rm yr}$.
Figure 3: Measurement precision (upper) of the spin parameter as functions of the orbital inclination angle $i_{2}$ of the second pulsar. Two dashed lines show the measurement precision from only one pulsar. The blue curve in the upper panel shows the combined result. The lower panel shows the improvement factor over the inner pulsar. Predictions from Eq. (1) are denoted with vertical lines.
@@
The simulation results are consistent with the prediction from the analysis we presented before. From Eq. (1), it is obvious that the shape of the hyperbola of each pulsar is controlled by their inclination angles, which determine the leading-order degeneracy. Moreover, from Eq. (1) one can also predict the position of the three degeneracy peaks for the combined results in Fig. 3, where the improvement from combining a second pulsar is the smallest. The central peak at $i=90^{\circ}$ originates from the denominator of the first term in Eq. (1), while the other two peaks are attributed to the situation where two hyperbolas are tangent to each other at the point they intersect. We show the predicted peak positions from Eq. (1) in Fig. 3 with vertical lines.
@@
In this work we have ignored the complex environments near the Sgr A*, which will complicate the measurement Merritt et al. (2010); Liu et al. (2012); Psaltis et al. (2016); Hu et al. (2023). In general, pulsars with larger orbital periods will suffer from more perturbations caused by the mass distributions around Sgr A*, which may even spoil the measurement. It is suggested that one may only use the timing data around the periastron passages and treat each periastron passage incoherently Psaltis et al. (2016). This will enlarge the measurement uncertainty of the spin from timing a single pulsar, where second-order time derivatives are required but hard to measure during the short periastron passage. However, as only the leading-order time derivatives are required in the method we proposed here, it is expected to have a good measurement even only the timing data around the periastron passages are used. On the other hand, by combing the timing observations of more pulsars, one will have the ability to model the mass distribution around Sgr A* that will be of vastly useful for studies of the GC region.

### v2

@@ Abstract
Measuring the BH spin is crucial for gravity tests. Determining the spin of Sgr A* based on the timing of a single pulsar suffers from the so-called leading-order degenracy among spin parameters Liu et al. (2012); Zhang and Saha (2017). The measurable secular changes of the pulsar orbit in a short time, caused by the frame-dragging effect Barker and O’Connell (1975), are the linear rate of the periastron precession, $\dot{\omega}$, and the linear change of the projected semimajor axis of the pulsar orbit, $\dot{x}$. However, these two quantities cannot fully determine the BH spin which has three components. To break the degeneracy, one needs at least one accurate measurement of higher-order time derivatives such as $\ddot{\omega}$ or $\ddot{x}$ Liu et al. (2012), which in general requires a much longer observation time span or a pulsar in an orbit with very short orbital period whose relativistic effects are large.
There are still two possible ways to break the leading-order degeneracy and provide a better spin determination without requiring a pulsar in a rare orbital situation. One is to combine the measurement of the proper motion of the pulsar from astrometric observations Zhang and Saha (2017). One can infer the secular change in the longitude of the ascending node, $\dot{\Omega}$, from the proper motion of the pulsar and provide an additional constraint on the spin of the central BH. However, an expected astrometric accuracy in the order of $10\,\mu{\rm as}$ Fomalont and Reid (2004) is still not enough and contributes little when combining with the timing data Zhang and Saha (2017). The other possibility is to combine the observation of another pulsar with a different orbital inclination $i$, as the leading-order degeneracy mainly depends on $i$ Zhang and Saha (2017). Considering a potentially large pulsar population in the GC, finding two or more pulsars orbiting around Sgr A* in the future may provide us a more likely way in measuring the spin of Sgr A*.
@@
Combining two pulsars.—When the changes in the Keplerian parameters of a pulsar orbit are slow over the observation time span, the direct way to characterize secular effects is to fit the timing data with timing parameters varying linearly in time, such as $\omega=\omega_{0}+\dot{\omega}\,t$ and $x=x_{0}+\dot{x}\,t$ Damour and Taylor (1992). Different from the binary pulsar systems that are regularly timed at present, in which even for the very precisely timed pulsars only some of them have measurents of $\dot{\omega}$ Chakrabarti et al. (2021); Donlon et al. (2024), in a pulsar-SMBH system, the leading-order drivatives will be measured soon after the start of observation and have a rather high measurement precision due to the large relativistic effects from the SMBH Liu et al. (2012); Psaltis et al. (2016); Zhang and Saha (2017); Hu et al. (2023). Here we use the two combinations introduced by Liu et al. (2012) for the leading-order derivatives and other timing observables, which are $\mathcal{X}\equiv-\dot{x}s_{i}^{2}\big{(}x\hat{\Omega}\big{)}^{-1}$ and $\mathcal{W}\equiv(\dot{\omega}-\dot{\omega}_{M})s_{i}^{2}\hat{\Omega}^{-1}$, where $s_{i}=\sin i$, $\hat{\Omega}=4\pi\beta_{O}^{3}/P_{b}(1-e^{2})^{3/2}$ and $\dot{\omega}_{M}=6\pi\beta_{O}^{2}/P_{b}(1-e^{2})$, with $\beta_{O}=\big{(}2\pi GM/c^{3}P_{b}\big{)}^{1/3}$ the orbital velocity parameter introduced by Damour and Taylor (1992). Note that $s_{i}$ and $M$ in these combinations can be measured precisely from the Shapiro delay Shapiro (1964) that is large enough here even for face-on orbits Liu et al. (2012); Hu et al. (2023).
@@
To combine the timing results of two pulsars and determine the spin of the central BH, one has to translate the measurement constraints of both pulsars in a common parameter space. In this case, the natural choice is the $\chi$-$\chi_{\lambda}$ plane, where $\chi_{\lambda}$ is the projection of the dimensionless spin $\chi$ of the BH in the line of sight direction. After some tedious simplification, we show that, for each pulsar $k$ ($k=1,2$), timing observations give one constraint represented by a hyperbola,
$\frac{\mathcal{X}_{k}^{2}}{s_{i_{k}}^{2}(1-s_{i_{k}}^{2})}+\frac{\mathcal{W}_{k}^{2}}{s_{i_{k}}^{2}(1+3s_{i_{k}}^{2})}=\chi^{2}-\frac{1+3s_{i_{k}}^{2}}{(1-3s_{i_{k}}^{2})^{2}}\left(\chi_{k}+\frac{3\mathcal{W}_{k}c_{i_{k}}}{1+3s_{i_{k}}^{2}}\right)^{2}\,,$ (1)
@@
where $c_{i_{k}}=\cos{i_{k}}$. If one observes two or more pulsars, all these curves in the $\chi$-$\chi_{\lambda}$ plane should intersect at one point that gives $\chi$ and $\chi_{\lambda}$.
@@
Figure 1: Illustration of the determination of the spin of Sgr A* with measurements of two pulsars in the $\chi$-$\chi_{\lambda}$ plane. The square symbol located at $(0.36,0.6)$ marks the assumed true value of the BH spin. The solid curves show the constraints from two pulsars, while the dashed lines correspond to the $i\leftrightarrow\pi-i$ ambiguity. Shaded region is excluded by the condition $\left|\chi_{\lambda}\right|\leq\chi$.
@@
Figure 1 shows an illustration of using the timing of two pulsars to determine the spin of Sgr A* in the $\chi$-$\chi_{\lambda}$ plane. The system parameters are randomly chosen to avoid the accidental degeneracy which will be discussed further in the numerical simulation section. For each pulsar $\alpha$, the measurements of $({\cal X}_{k},{\cal W}_{k})$ give a hyperbola shown by the solid curve, and the $i_{k}\leftrightarrow\pi-i_{k}$ ambiguity gives another hyperbola shown by the dashed line. In general, there can exist at most eight different solutions for $(\chi,\chi_{\lambda})$. A rough measurement of any second-order derivatives, for example, the $\ddot{\omega}$ for one of the pulsar, as shown by the vertical red line, or timing of a third pulsar can reduce this degeneracy into two solutions whose $\chi$ has the same value and $\chi_{\lambda}$ has opposite signs.
@@
The methodology described above can be easily extended for combining more pulsars. For each additional pulsar, it provides an independent test of GR as all these curves should intersect at one point within their measurement uncertainties. On the other hand, if one assumes that GR is correct, timing more pulsars can be a unique probe for us to explore and model the (stellar and gas) mass distribution near the GC Merritt et al. (2010); Liu et al. (2012); Hu et al. (2023); Carleo and Ben-Salem (2023). As only the leading-order time derivatives and a rough value of a higher-derivative parameter are used, one is expected to obtain a high-precision measurement after a short time observing two pulsars.
@@
The full parameter set, $\bm{\Theta}=\bm{\Theta}_{\rm BH}\cup\bm{\Theta}_{\rm PSR}$, in the timing model for a single pulsar includes the parameters of the SMBH, $\bm{\Theta}_{\rm BH}=\{M,\chi,q,\lambda,\eta\}$, where $\lambda$ and $\eta$ describe the direction of the spin, and the parameters of the pulsar, $\bm{\Theta}_{\rm PSR}=\{P_{b},e,f,i,\omega,N_{0},\nu,\dot{\nu}\}$, where $f$ is the initial orbital phase and $\{N_{0},\nu,\dot{\nu}\}$ relate the pulsar’s rotation number $N$ and pulsar’s proper time $T$ via $N(T)=N_{0}+\nu T+\dot{\nu}T^{2}/2$. We have set the longitude of the ascending node of the pulsar orbit, $\Omega$, to be zero as it is not an observable when ignoring the proper motion Kopeikin (1994). When combining the timing of two or more pulsars, one should have one set of $\Theta_{\rm PSR}$ for each pulsar while $\Theta_{\rm BH}$ is common for different pulsars. In addition, the difference between the longitudes of the ascending node, $\Delta\Omega=\Omega_{2}-\Omega_{1}$, should also be included in the timing model Ransom et al. (2014). As we can have an overall rotation of the system around the line of sight direction, we fix $\Omega_{1}=0$ in the following discussions and we denote the pulsar with a smaller orbital period as PSR1 or the inner pulsar.
We use the Fisher matrix method for parameter estimation Damour and Deruelle (1986). The log-likelihood function $\mathcal{L}$ in the covariance matrix $C_{\mu\nu}=\big{(}\partial^{2}\mathcal{L}/\partial\Theta^{\mu}\partial\Theta^{\nu}\big{)}^{-1}$ is simply the sum from two pulsars, $\mathcal{L}=\mathcal{L}_{1}+\mathcal{L}_{2}$, and
$\mathcal{L}_{k}=\frac{1}{2\nu_{k}^{2}}\sum_{a=1}^{N^{(k)}_{\rm TOA}}\frac{\left[N^{(k)}_{a}(\bm{\Theta})-N^{(k)}_{a}(\bar{\bm{\Theta}})\right]^{2}}{\left(\sigma^{(k)}_{\rm TOA}\right)^{2}}\,,\ \ \ k=1,2\,$ (2)
where $N^{(k)}_{a}(\Theta)=N\Big{(}\Theta;t_{a}^{(k)\,{\rm TOA}}\Big{)}$ is the pulsar rotation number calculated by the timing model corresponding to the $a$-th pulse arrived at $t_{a}^{(k)\,{\rm TOA}}$. $\bar{\bm{\Theta}}$ denotes the true system parameters and $\sigma_{\rm TOA}^{(k)}$ is the timing precision for pulsar $k$. As we do not consider the interaction between the two pulsars, the properties of the Fisher matrix method and the timing model allow one to construct the full Fisher matrix $F\equiv C^{-1}$ though the combination of the Fisher matrices $F_{i}$ of each single pulsar. This simplification is important for combining more pulsars.
@@
We assume a timing precision $\sigma_{\rm TOA}=1\,{\rm ms}$ and an observation time span $T_{\rm obs}=5\,{\rm yr}$ for both pulsars, which are realistic for future instruments such as the Square Kilometre Array (SKA) and the next-generation Very Large Array (ngVLA) Weltman et al. (2020); Bower et al. (2018). In the upper panel of Fig. 2 we show the measurement precision of the spin of Sgr A* from combining the timing of two pulsars (solid curves) compared to the results obtained from the inner pulsar only (dashed line). Here we fix the orbital period of the inner pulsar to be $P_{b1}=2.3\,{\rm yr}$ and orbital eccentricity $e_{1}=0.8$ while varying the orbital period $P_{b2}$ and eccentricity $e_{2}$ of the second pulsar. The lower panel of this figure shows the improvement factor $F_{\rm imp}=\delta\chi_{\rm single}/\delta\chi_{\rm combine}$. For this case, the inner pulsar with an orbital period $P_{b1}=2.3\,{\rm yr}$ can only constrain $\delta\chi/\chi$ to $\sim 15\%$. Combining a second pulsar even with a large orbital period can provide an improvement factor of $\sim 10$ because of the breaking of the leading-order degeneracy. The improvement factor from combining a second pulsar will be smaller if the inner pulsar has a tighter orbit for which itself can break the leading-order degeneracy. Another case study of $P_{b1}=0.5\,{\rm yr}$ shows an improvement factor of $2\sim 3$ when combined with a second pulsar whose orbital period $P_{b2}\lesssim 3\,{\rm yr}$.
Figure 3: Measurement precision (upper) of the spin parameter as functions of the orbital inclination angle $i_{2}$ of the second pulsar. Two dashed lines show the measurement precision from only one pulsar. The blue curve in the upper panel shows the combined result. The lower panel shows the improvement factor over the inner pulsar. The predicted positions of the degeneracy peaks from Eq. (1) are denoted with vertical lines.
@@
The simulation results are consistent with the prediction from the analysis of Eq. (1) as shown below. From Eq. (1), it is obvious that the shape of the hyperbola of each pulsar is controlled by their inclination angles, which determine the leading-order degeneracy. Moreover, from Eq. (1) one can also predict the position of the three degeneracy peaks for the combined results in Fig. 3, where the improvement from combining a second pulsar is the smallest. The central peak at $i=90^{\circ}$ originates from the denominator of the first term in Eq. (1), while the other two peaks are attributed to the situation where two hyperbolas are tangent to each other at the point they intersect. We show the predicted peak positions from Eq. (1) in Fig. 3 with vertical lines. Note that $s=0$ and $s=1/\sqrt{3}$ do not lead to additional degeneracy peaks as they behave normally when considering error propagation in Eq. (1).
@@
In this work we have ignored the complex environments near the Sgr A*, which will complicate the measurement Merritt et al. (2010); Liu et al. (2012); Psaltis et al. (2016); Hu et al. (2023). In general, pulsars with larger orbital periods will suffer from more perturbations caused by the mass distributions around Sgr A*, which may even spoil the measurement. Estimations made by previous studies suggest that for pulsars with orbital period $P_{b}\lesssim 3\,{\rm yr}$, the orbital advance rate caused by the SMBH spin is still larger than that caused by the stellar perturbation, which indicates the measurability of the SMBH spin with pulsar-SMBH systems Liu et al. (2012). It is also suggested that one may only use the timing data around the periastron passages and treat each periastron passage incoherently Psaltis et al. (2016). This will enlarge the measurement uncertainty of the spin from timing a single pulsar, where second-order time derivatives are required but hard to measure during the short periastron passage. However, as only the leading-order time derivatives are required in the method we proposed here, it is expected to have a good measurement even only the timing data around the periastron passages are used. On the other hand, by combing the timing observations of more pulsars, one will have the ability to model the mass distribution around Sgr A* that will be of vastly useful for studies of the GC region.